# The Role of Entanglement in Quantum Reservoir Computing with Coupled Kerr Nonlinear Oscillators

### v2

@@ Acknowledgements.
Appendix A Entanglement Advantage
To further illustrate the entanglement advantage discussed in the main text, we provide additional figures and analysis.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ Appendix A Entanglement Advantage
Fig. 9 shows the full binned relation between NRMSE and entanglement across multiple dissipation rates. The trend confirms that for sufficient dissipation, moderate entanglement is associated with better prediction performance, up to a saturation point. This supports our claim that optimal entanglement lies in an intermediate regime.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ C.1 Entanglement Advantage
Notably, we consolidated the results from the Figures 12, 13, and 14, and performed a binned analysis to produce Fig. 17 which demonstrates that the relationship between entanglement and predictive performance remains consistent with the linear task results: moderate entanglement corresponds to reduced average and worst-case errors. This confirms that the observed entanglement advantage is not an artifact of the specific linear task used in the main text but persists in a more complex and nonlinear forecasting benchmark.
@@
Figure 12: Same as Fig. 3 but for the NARMA task.
@@ C.2 Dissipation & Dephasing
Refer to caption Refer to caption Refer to caption

### v3

@@ Acknowledgements.
Appendix A Effects of Dissipation and Dephasing on the Overall Entanglement-Performance Relationship
To further illuminate the relationship between entanglement and performance discussed in the main text, we provide additional figures and analysis.
Refer to caption Refer to caption
Refer to caption Refer to caption
@@ Appendix A Effects of Dissipation and Dephasing on the Overall Entanglement-Performance Relationship
Fig. 9 shows the full binned relation between NRMSE and entanglement across multiple dissipation rates. The trend confirms that for sufficient dissipation, moderate entanglement is associated with better prediction performance, up to a saturation point. This supports our claim that optimal entanglement lies in an intermediate but non-zero regime.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
@@ C.1 Entanglement-Performance Relationship
Notably, we consolidated the results from the Figures 12, 13, and 14, and performed a binned analysis to produce Fig. 15 which demonstrates that the relationship between entanglement and predictive performance remains consistent with the linear task results: moderate but non-zero entanglement corresponds to minimum errors. This confirms that the observed entanglement-performance relationship is not an artifact of the specific linear task used in the main text but persists in a more complex and nonlinear forecasting benchmark.
@@
Figure 12: Same as Fig. 3 for the NARMA task.